OpenStudy (anonymous):

Help with mathematical induction? The problem is (3n-2)^2 = (n(6n^2 -3n -1))/2 I know that n=1 is okay. Both sides equal 1. But when I start plugging in k+1, I get lost. Can someone help me test k+1?

OpenStudy (zehanz):

You are proving it with Induction. This means you have to do the following: 1. Check to see if it is true for n=1. Done. 2. Suppose it is true for n=k, then prove it's true for n=k+1. Just to be safe, 2. means:\[(3k+2)^2=\frac{ k(6k^2-3k-1) }{ 2 }\]is true. Not let's plug in n=k+1. Left side:\[(3(k+1)+2)^2\] Right side:\[\frac{ (k+1)(6(k+1)^2-3(k+1)-1) }{ 2 }\] I suppose you have come this far.

OpenStudy (campbell_st):

ok.... you need to find the term in (k + 1) \[T_{k +1} = (3(k + 1) -2)^2 \] and the sum of k terms is \[S_{k} = \frac{k(6k^2 -2k -1)}{2}\] you need to show \[T_{k + 1} + S_{k} = \frac{(k + 1)(6(k + 1)^2 - 3(k + 1) - 1)}{2}\]

OpenStudy (campbell_st):

so looking at \[T_{k + 1}+S_{k} = \frac{ 2(3k +1)^2}{2} + \frac{k(6k^2 -3k -1)}{2} \]

OpenStudy (anonymous):

@ZeHanz, thanks. Yes, I have gotten that far. It is what I have to do from there that I am confused. @campbell_st , thanks. I think I understand what you mean so far.

OpenStudy (campbell_st):

which I think is \[T_{k +1}+ S_{k} = \frac{ 6k^3 + 15k^2 + 11k + 2}{2}\]

OpenStudy (campbell_st):

now its expanding \[S_{k + 1}\] to show its equal

OpenStudy (zehanz):

Look at campbell_st's explanation...

OpenStudy (campbell_st):

the first part of the expansion is \[S_{k + 1} = \frac{(k + 1)(6(k^2 + 2k + 1) -3k - 3 - 1)}{2}\] which then becomes \[S_{k + 1} = \frac{(k + 1)(6k^2 + 9k + 2)}{2}\] it shoule be straight forward from here to prove for k + 1 I've done it and I know it works..

OpenStudy (campbell_st):

all you need to do is show \[T_{k + 1} + S_{k} = S_{k + 1}\]

OpenStudy (anonymous):

I see what you're doing and I'm trying to replicate it. I can't seem to find how tk+1 + sk = sk+1. But you said it was true right?

OpenStudy (anonymous):

@campbell_st, is this statement false?

OpenStudy (campbell_st):

the statement is true this is what you are trying to prove \[T_{k + 1} + S_{k} = S_{k + 1} \] so if you know the sum to a specific term and then add the next term you should get the sum to the k + 1 th term so \[\frac{6k^3 + 15k^2 + 11k + 2}{2} = \frac{(k + 1)(6k^2 + 9k + 2)}{2} \] distribute on the right hand side gives \[\frac{6k^3 + 15k^2 + 11k + 2}{2} = \frac{ 6k^3 + 9k^2 + 2k + 6k^2 + 9k + 2}{2}\] collect the like terms on the right and you'll see that boths sides are equal, as required.

OpenStudy (campbell_st):

hope this make sense.

OpenStudy (anonymous):

OHHHHH I see. Thanks so much man. I actually understand this now. :P